# Who Does the Giant Number Pile Like Best: Analyzing Fairness in Hiring Contexts

### v1

@@ Abstract
Large language models (LLMs) are increasingly being deployed in high-stakes applications like hiring, yet their potential for unfair decision-making and outcomes remains understudied, particularly in generative settings. In this work, we examine the fairness of LLM-based hiring systems through two real-world tasks: resume summarization and retrieval. By constructing a synthetic resume dataset and curating job postings, we investigate whether model behavior differs across demographic groups and is sensitive to demographic perturbations. Our findings reveal that race-based differences appear in approximately 10% of generated summaries, while gender-based differences occur in only 1%. In the retrieval setting, all evaluated models display non-uniform selection patterns across demographic groups and exhibit high sensitivity to both gender and race-based perturbations. Surprisingly, retrieval models demonstrate comparable sensitivity to non-demographic changes, suggesting that fairness issues may stem, in part, from general brittleness issues. Overall, our results indicate that LLM-based hiring systems, especially at the retrieval stage, can exhibit notable biases that lead to discriminatory outcomes in real-world contexts.
Who Does the Giant Number Pile Like Best: Analyzing Fairness in Hiring Contexts
Preethi Seshadri UC Irvine preethis@uci.edu                        Seraphina Goldfarb-Tarrant Cohere seraphina@cohere.com
@@ 1 Introduction
Large language models (LLMs) are increasingly being used for real-world, high-stakes domains, including medicine, law, and hiring. As LLMs are incorporated into critical decision-making processes, ensuring reliable and responsible usage is paramount. A key aspect of this effort includes anticipating and preventing specific risks and harms, such as identifying biases that could unfairly disadvantage certain demographic groups in hiring screening processes.
@@
Prior work focuses on two types of harms: representational harms (i.e., representing certain groups in a less favorable light than others, demeaning them, or failing to recognize their existence altogether) and allocational harms (i.e., allocating resources or opportunities or unfairly to different social groups) (Barocas et al., 2017; Blodgett et al., 2020). There is a substantial body of work that analyzes representational harms in LLMs (Zhao et al., 2018; Abid et al., 2021; Kirk et al., 2021; Cheng et al., 2023; Gadiraju et al., 2023), yet allocational harms—which are the primary harm at play in high-stakes situations—remain understudied. Given the growing use of LLMs in consequential applications, it is especially important to examine whether models produce unfair outcomes or performance gaps between different groups (allocational harms, also referred to as allocational fairness).
Although there are a few studies on LLM harms in real-world settings (e.g., hiring) (Tamkin et al., 2023; An et al., 2024; Haim et al., 2024; Nghiem et al., 2024), there is minimal work closely simulating actual deployment scenarios. By using highly simplified setups and tasks, these prior works may fail to predict real-world outcomes and as a result provide an incomplete assessment of harms. Investigations into LLM harms should ensure ecological validity (Goldfarb-Tarrant et al., 2021; Cao et al., 2022); they should either be grounded in realistic scenarios and tasks that match how these systems are used in practice, or use a proxy that has been proven to be predictive of real world outcomes.
@@
Additionally, these existing studies have cast their investigations of allocational harms in LLMs as discrete classification tasks (such as binary accept/reject decisions) or quantitative predictions (such as determining salary levels). This makes the implementation and evaluation more straightforward, but does not match the real use of generative LLMs (Kelly, 2023; Yin et al., 2024). In contrast, there is limited work on allocational harms in generative settings (Wan et al., 2023), like summarization, since measuring how generated text might create disparities is more open-ended and complex than analyzing classification predictions.
In this work, we examine whether LLMs are fair in real-world contexts for hiring. We choose hiring due to the growing adoption of LLMs in hiring tools (Sánchez-Monedero et al., 2020). Hiring is a high-stakes domain with clear societal risks – unfair systems may discriminate against already marginalized groups, risking further underrepresentation throughout the hiring pipeline (Dastin, 2018; Raghavan et al., 2020; Suresh and Guttag, 2021). We focus on two tasks that mirror how LLMs are integrated into hiring and recruitment workflows: resume summarization and resume retrieval with respect to a job post (Herman, 2024; Humanly, 2024). We then examine whether model behavior in these tasks is sensitive to gender and race demographic perturbations in resumes (e.g., name perturbations or tailored extracurricular content). Specifically, we ask the following questions:
  • RQ1: Do model generated summaries differ meaningfully across demographic groups?
  • RQ2: In a retrieval setting, do models disparately select resumes across demographic groups?
  • RQ3: In a retrieval setting, how sensitive are model rankings to demographic and non-demographic perturbations in resumes?
In this work, we: (1) create a synthetic resume dataset and curate a job post dataset, (2) propose metrics to measure fairness in generative and retrieval settings for two real-world hiring tasks, validated by an expert human preference study, and (3) perform a study investigating the fairness of an automated LLM hiring system based on real world usage, for 7 models (3 generative, 4 retrieval). [^1] The data and code can be found at https://github.com/preethisesh/hiring_fairness.
@@
In the summarization setting, on average, we observe meaningful differences between summaries generated for different racial groups nearly 10% of the time, as opposed to only 1% for gender (RQ1). On the other hand, for retrieval, we find that all models are prone to non-uniformly selecting resumes across demographic groups (RQ2). In addition, all models produce rankings that are highly sensitive to both name and race demographic perturbations (RQ3). We also find that model sensitivity to non-demographic perturbations is notable and sometimes even on par with demographic perturbations (RQ3), suggesting that fairness issues may largely be due to model brittleness as a whole, rather than encoded biases only. Overall, our results demonstrate that an LLM hiring system with automated resume retrieval and summarization exhibits considerable bias, primarily stemming from the initial retrieval stage.
@@ 2 Methodology
To study fairness in hiring, we consider an LLM-based pipeline with two components: resume retrieval with respect to a job post (using an embedding model), and resume summarization (using an LLM). [^2] We analyze both components because they reflect real-world usage of automation to streamline hiring processes. We focus on summarization first because it is more neglected in research, though in a pipeline it would come after retrieval (as summarization would be of retrieved resumes). We investigate the impact of perturbing the demographic group of a resume on each component of the pipeline and look for: (1) systematic differences in generated resume summaries and (2) changes in similarity of a resume to a job posting, and as a result its ranking in a resume set.
Let $D$ represent a set of resumes, where each resume $d\in D$ has demographic label $l(d)$. For each resume, let $d^{\prime}$ represent any perturbed version of $d$ where $l(d^{\prime})\neq l(d)$. For each $d$, there can be several perturbed versions $d^{\prime}$ along various demographic dimensions (e.g., gender or race, to name a few). The resume content remains largely unchanged except for the demographic perturbation.
@@ 2.1 Summarization
Perturbed resumes are by design highly similar to original resumes, so we expect generated summaries for original and perturbed resumes to also maintain a high degree of similarity. In other words, we are testing for invariance, since we expect the output to change minimally after perturbing the input (Ribeiro et al., 2020).
In order to assess invariance, we need a way to measure whether the original and perturbed summaries differ meaningfully. For cost and scalability reasons, we rely on an automated approach—human preferences are expensive and cannot be collected frequently, especially if these preferences are sourced from HR staff. Additionally, collecting human preferences for a specific dataset does not extend to other analyses with new datasets.
Automated measures can evaluate whether specific properties of generated summaries differ, where differences in these properties would be undesirable in practice. For instance, summaries for a specific demographic group should not be written more positively than summaries for other demographic groups; this could lead to disparity in hiring outcomes. We use the following measures as proxies for undesirable variation that could influence the decision of an HR staff reading the summary: reading ease, reading time, polarity, subjectivity, and regard (Appendix A.5). We verify that these measures capture meaningful differences in practice by conducting a preference task annotated by HR staff. This study shows that these are good proxies for human preferences (Appendix A.6).
@@
We compute an invariance metric that aggregates the proposed measures for both original and perturbed resume summaries. For each measure, we do a paired t-test between scores for original and perturbed summaries, repeated for each demographic perturbation. We then calculate how often the null hypothesis (that the mean difference between paired summaries is zero) is rejected, i.e., how often invariance is violated. We choose a significance level of $\alpha=0.05$, and apply Bonferroni correction to account for Type 1 errors with multiple comparisons (Bland and Altman, 1995).
@@ 2.2 Retrieval
In contrast to summarization, which relies on only the resume, retrieval utilizes both a query (i.e., job posting) and candidates to select (i.e., resumes). The top-$n$ of the resulting resumes sorted by similarity then make it to the next stage of the pipeline. Similar to the previous setting, we assume that demographically perturbing a resume should have no impact on its relevance to a job posting.
Given an embedding model $\mathcal{M}$ and a similarity measure $sim$, we compute the similarity between embeddings for resume $d$ and a job posting $p$. In practice, we use cosine similarity (also used in Wilson and Caliskan (2024)). We define the set $S(p)$, which represents the set of similarity values between each resume $d\in D$ and job posting $p$:
$S(p)=\{sim(\mathcal{M}(d),\mathcal{M}(p))\mid d\in D\}$
We transform the similarity values between each resume and job posting, $s_{i}$, into a rank such that lower ranks indicate higher similarity (note: the highest similarity value receives rank 1):
$\text{rank}_{p}(s_{i})=|\{s_{j}\in S(p)\mid s_{j}>s_{i}\}|+1$
Let $D_{n}(p)$ represent the set of top-$n$ resumes from $D$, which are the resumes with the $n$ lowest ranks (i.e., highest similarities) for job posting $p$:
$D_{n}(p)=\{d_{i}\in D\mid\text{rank}(s_{i})\leq n\}$
@@
We compute two fairness metrics for retrieval, non-uniformity, proposed by Wilson and Caliskan (2024), and exclusion, which we introduce below. Further intuition about these metrics are provided in Appendix A.7.
@@
evaluates whether the top-$x$(where $x$ is a percentage instead of a fixed number $n$) of resumes are uniformly distributed with respect to the demographic groups being considered. To compute non-uniformity, we consider the set of top-$x$of resumes retrieved from the combined pool of all demographically perturbed versions, which we denote as $D_{x}^{\prime}(p)$. Following the approach from Wilson and Caliskan (2024), we perform a chi-squared goodness-of-fit test to determine whether the demographic makeup of $D_{x}^{\prime}(p)$ deviates from the uniform distribution.
@@
evaluates how often resumes in the set of top-$n$ resumes are excluded (i.e., the ranking falls outside top-$n$) after perturbation. Ideally, $\mathcal{M}$ should be robust to demographic perturbations, yielding nearly identical similarity scores and rankings for both resume $d$ and demographically perturbed version $d^{\prime}$. Exclusion directly tests this behavior by measuring the extent to which the set of top-$n$ resumes differs after perturbation.
@@ 3 Experimental Setup
In this section, we provide an overview of the datasets, perturbations, and models used in our evaluation.
@@ 3.1 Data
Resumes are sourced through social media platforms including LinkedIn, Slack, and X. Given the authors’ professional circles, the sample skews heavily toward tech and academic professionals. For privacy reasons, we anonymize resumes using Presidio anonymize, [^3] https://microsoft.github.io/presidio/anonymizer/ which masks PII entities. To further mitigate privacy concerns and enable us to release the dataset, we use the collected resumes as examples to generate synthetic resumes, instead of using original resumes in our analysis. We generate 525 resumes across 21 professions using Cohere’s Command-R model (Cohere, 2024). All synthetic resumes are free of explicit demographic information, until we add it as part of experimentation.
In addition to our curated resume set, we use a publicly available resume dataset from Kaggle (Bhawal, 2021). [^4] Has a CC0: Public Domain License These resumes differ in two notable ways: (1) the Kaggle resumes are less structured and formatted than synthetically generated resumes, and (2) they include a more diverse set of fields (e.g., apparel, construction, fitness, etc.). We sample 1175 Kaggle resumes across 24 fields.
@@
Our resume dataset consists of two types: realistic synthetic resumes generated for specific roles (e.g., data analyst) and actual resumes from Kaggle labeled with broader field categories (e.g., construction). For each profession/field (we choose 11 from generated resumes and 11 from Kaggle resumes), we carefully select 7 detailed LinkedIn job postings that clearly describe the role and outline qualifications. For Kaggle resumes, we match postings to roles within each broader field (e.g., choosing “construction manager” postings for the construction field). In total, we collect 154 job postings across different professions.
@@ 3.2 Demographic Perturbations
We use names as a proxy for gender and racial information. Initially, all resumes are free of names. We then add names using the curated set of names from Yin et al. (2024). [^5] Uses voter registration data from North Carolina to identify demographically-distinct names. We consider four demographic groups, each with 100 unique names: Black female (FB), White female (FW), Black male (MB), and White male (MW). Following Wilson and Caliskan (2024), we use the demographically neutral “Williams” as the last name for all groups and only vary first name.
In actual resumes, demographic information can be encoded in more than just the name. Therefore, for synthetically generated resumes, we perform an additional augmentation step that adds extracurricular information to the resumes (similar to Glazko et al. (2024)). We use Command-R to augment resumes by including awards, clubs and leadership, and mentorship and volunteering experiences that are reflective of the individual’s background and identity (more details in Appendix A.3 and A.4). In addition to explicit gender and race information from names, extracurricular information can reinforce demographic signal by providing further explicit and implicit cues.
(a) GPT-4o
(b) Command-R
(c) Mixtral 8x7b
Figure 1: Summarization Results: Percentage of t-tests for which we reject null hypothesis (i.e., invariance is violated), separated by completion model and perturbation.

### v2

# Small Changes, Large Consequences: Analyzing the Allocational Fairness of LLMs in Hiring Contexts

@@ Abstract
Large language models (LLMs) are increasingly being deployed in high-stakes applications like hiring, yet their potential for unfair decision-making remains understudied in generative and retrieval settings. In this work, we examine the allocational fairness of LLM-based hiring systems through two tasks that reflect actual HR usage: resume summarization and applicant ranking. By constructing a synthetic resume dataset with controlled perturbations and curating job postings, we investigate whether model behavior differs across demographic groups. Our findings reveal that generated summaries exhibit meaningful differences more frequently for race than for gender perturbations. Models also display non-uniform retrieval selection patterns across demographic groups and exhibit high ranking sensitivity to both gender and race perturbations. Surprisingly, retrieval models can show comparable sensitivity to both demographic and non-demographic changes, suggesting that fairness issues may stem from broader model brittleness. Overall, our results indicate that LLM-based hiring systems, especially in the retrieval stage, can exhibit notable biases that lead to discriminatory outcomes in real-world contexts.
Small Changes, Large Consequences:
Analyzing the Allocational Fairness of LLMs in Hiring Contexts
Preethi Seshadri UC Irvine preethis@uci.edu          Hongyu Chen Cohere charlie@cohere.com          Sameer Singh UC Irvine sameer@uci.edu          Seraphina Goldfarb-Tarrant Cohere seraphina@cohere.com
@@ 1 Introduction
Large language models (LLMs) are increasingly being adopted in real-world, high-stakes domains such as hiring (Boston Consulting Group, 2025), where they assist HR teams with tasks like resume screening and candidate matching. As LLMs are incorporated into critical decision-making processes, ensuring fair and responsible deployment is essential, especially when the outcomes can profoundly impact individuals’ career prospects (Dastin, 2018; Raghavan et al., 2020; Sánchez-Monedero et al., 2020; Suresh and Guttag, 2021).
@@
A key aspect of developing responsible LLM systems includes anticipating and preventing specific risks and harms, such as allocational harms (i.e., allocating resources or opportunities unfairly to different social groups, also called allocational fairness) (Barocas et al., 2017; Blodgett et al., 2020). This is especially important in automated hiring pipelines, since models may produce unfair outcomes and reinforce systemic inequalities. While there is a substantial body of work that analyzes representational harms (i.e., representing certain social groups negatively, demeaning them, or erasing their existence) in LLMs (Zhao et al., 2018; Abid et al., 2021; Kirk et al., 2021; Cheng et al., 2023; Gadiraju et al., 2023), allocational harms—which are the primary harm at play in high-stakes situations—remain understudied beyond discriminative systems.
@@
The few studies that evaluate allocational harms of LLMs (Tamkin et al., 2023; An et al., 2024; Haim et al., 2024; Nghiem et al., 2024) have primarily cast their investigations as discrete classification tasks (e.g., yes/no decisions) or quantitative predictions (e.g., determining salary levels), which do not capture how LLMs are deployed in applications like hiring (Kelly, 2023). As a result, these highly simplified setups may inadequately predict real-world outcomes and assess harms. Investigations of LLM harms must ensure ecological validity (Blodgett et al., 2021; Goldfarb-Tarrant et al., 2021; Cao et al., 2022); they should be grounded in realistic scenarios and tasks that match how these systems are used in practice, or use a proxy that is predictive of real world outcomes. Yet there is limited work on allocational harms in generative settings without adding a simplification layer, with Wan et al. (2023) being a notable exception, since measuring how generated text might create disparities is more open-ended and complex than analyzing classification predictions.
In this work, we examine whether LLMs behave fairly in real-world hiring contexts. We focus on two critical tasks that mirror how LLMs are integrated into hiring workflows (Herman, 2024; Humanly, 2024): (1) ranking candidates with respect to a job posting and (2) summarizing resumes, as illustrated in Figure 1. These tasks represent key stages where automation can influence which candidates are surfaced and considered for a role. To evaluate fairness, we examine whether models are sensitive to gender and race perturbations in resumes. We investigate the following questions:
  • RQ1: Do generated summaries differ meaningfully across demographic groups? [^1] We study summarization first, since it is less explored from an allocational harms perspective.
  • RQ2: Do models disparately select resumes across demographic groups?
  • RQ3: How sensitive are model rankings to demographic and non-demographic perturbations in resumes?
Figure 1: We investigate the fairness of an LLM hiring pipeline with a retrieval stage (ranks/filters the top-$n$ candidates with respect to a job post) and a summarization stage (generates resume summaries for filtered candidates). We assess fairness at each stage separately.11footnotemark: 1
To this end, we: (1) construct a new benchmark consisting of a synthetic resume dataset with controlled demographic perturbations (varying names and extracurricular content) and curated job postings, (2) design an evaluation framework with fairness metrics tailored to both generative and retrieval settings, validated by an expert human preference study, (3) conduct a comprehensive fairness analysis of 10 large language models (6 generative, 4 retrieval) based on real-world hiring tasks. We will make all data and code publicly available.
@@
Our results demonstrate that an LLM hiring system with automated resume retrieval and summarization exhibits considerable bias, primarily stemming from the retrieval stage. In the summarization setting, we observe meaningful differences in generated summaries up to 20% of the time between racial groups, compared to 3% for gender (RQ1). For retrieval, models non-uniformly select resumes across demographic groups up to 55% of the time (RQ2), and produce rankings that are highly sensitive to gender and race, with up to 74% of candidates being filtered out after demographic perturbation (RQ3). We also find that models exhibit high sensitivity to non-demographic changes, sometimes on par with demographic changes (RQ3), suggesting that fairness issues can stem from general model brittleness rather than demographic bias alone. Overall, our analysis reveals that even seemingly minor changes can lead to considerable disparities, raising concerns about the fairness and robustness of LLMs in hiring.
@@ 2 Methodology
To study fairness in hiring, we consider an LLM-based pipeline with two components: resume retrieval with respect to a job post (using an embedding model) and resume summarization (using an LLM). This pipeline is informed by interviews with several corporations that actively deploy LLMs for hiring; [^2] We cannot share details due to non-disclosure agreements. both components reflect real-world usage of automation to streamline hiring processes. We focus on summarization first because it is more neglected in research, though in a pipeline it would come after retrieval, as shown in Figure 1 (as summarization would be of retrieved resumes).
We propose two metrics, invariance violations (summarization) and exclusion (retrieval), to investigate allocational fairness in hiring. Specifically, these metrics quantify: (1) systematic differences in generated resume summaries [^3] Since recruiters may rely on summaries rather than full resumes, meaningful differences in summaries could directly affect hiring outcomes, leading to allocational harms. and (2) changes in the similarity of a resume to a job posting, and as a result its ranking in a resume set. We additionally benchmark the distribution metric of Wilson and Caliskan (2024) to study fairness in hiring, but note that their approach does not directly capture how perturbing a resume impacts resume screening outcomes, and includes only retrieval and not generative settings.
Let $D$ represent a set of resumes, where each resume $d\in D$ has demographic label $l(d)$. We denote $d^{\prime}$ as any perturbed version of $d$ where $l(d^{\prime})\neq l(d)$. For each $d$, there can be several demographically perturbed versions $d^{\prime}$ (e.g., gender or race perturbations). The resume content remains largely unchanged except for the demographic perturbation.
@@ 2.1 Summarization
Perturbed resumes are by design highly similar to original resumes, so we expect generated summaries for original and perturbed resumes to also maintain high similarity. In other words, we are testing for invariance; we expect the output to change minimally after perturbing the input (Ribeiro et al., 2020). To assess invariance, we need a way to measure whether the original and perturbed summaries differ meaningfully in the context of hiring. For cost and scalability reasons, we rely on an automated approach—human preferences are expensive and cannot be collected quickly enough to be used in model development, especially if these preferences are sourced from HR staff.
Automated measures can evaluate whether specific properties of generated summaries differ, in a way that could affect a human reader’s opinion. For instance, summaries for a specific demographic group should not be written more positively than summaries for other demographic groups; this could lead to disparities in hiring outcomes. We use the following measures as proxies for undesirable variation that could influence the decision of HR staff reading the summary: reading ease, reading time, polarity, subjectivity, and regard (Appendix A.8). We verify that these measures capture meaningful differences in practice by conducting a preference task annotated by HR staff. This study shows that these are good proxies for human preferences (Appendix A.9).
@@
For each measure, we perform a paired t-test between scores for original and perturbed summaries. We then calculate how often the null hypothesis (that the mean difference between paired summaries is 0) is rejected, i.e., how often invariance is violated. We choose a significance level of $\alpha=0.05$, and apply Benjamini-Hochberg correction to account for Type 1 errors with multiple comparisons (Benjamini and Hochberg, 1995).
@@ 2.2 Retrieval
In contrast to summarization, which relies only on the resume, retrieval uses both a query (job posting) and candidates to select (resumes). The top-$n$ of the resulting resumes sorted by similarity then make it to the next stage of the pipeline. We assume that demographically perturbing a resume should have minimal impact on its relevance to a job posting.
Given an embedding model $\mathcal{M}$ and a similarity measure simsim, we compute the similarity between embeddings for resume $d$ and job posting $p$. In practice, we use cosine similarity following Wilson and Caliskan (2024). We define the set $S(p)$, which represents the set of similarity values between each resume $d\in D$ and a given job posting $p$: $S(p)=\{sim(\mathcal{M}(d),\mathcal{M}(p))\mid d\in D\}$.
We transform the similarity values between each resume and job posting, $s_{i}$, into a rank such that lower ranks indicate higher similarity: $\text{rank}_{p}(s_{i})=|\{s_{j}\in S(p)\mid s_{j}>s_{i}\}|+1$. Let $D_{n}(p)$ represent the set of top-$n$ resumes from $D$, which are the resumes with the $n$ lowest ranks (i.e., highest similarities) for job posting $p$: $D_{n}(p)=\{d_{i}\in D\mid\text{rank}(s_{i})\leq n\}$.
@@
We compute two fairness metrics for retrieval, non-uniformity, proposed by Wilson and Caliskan (2024), and exclusion, which we introduce below. [^4] Further intuition and differentiation of these metrics are provided in Appendix A.11.
@@
assesses whether the top resumes are uniformly distributed across demographic groups. First, the set of top-xx% of resumes ($x$ being a percentage rather than a fixed number $n$) is retrieved from the combined pool of all demographically perturbed versions, denoted as $D_{x}^{\prime}(p)$. A chi-squared goodness-of-fit test is then used to check if the demographic composition of $D_{x}^{\prime}(p)$ deviates from the uniform distribution.
@@
evaluates how often resumes in the set of top-$n$ resumes are excluded (i.e., the ranking falls outside top-$n$) after perturbation. Ideally, $\mathcal{M}$ should be robust to demographic perturbations, yielding nearly identical similarity scores and rankings for both resume $d$ and demographically perturbed version $d^{\prime}$. Exclusion directly assesses allocational fairness by measuring how much the set of top-$n$ resumes differs after perturbation.
@@ 3 Experimental Setup
In this section, we describe the data, perturbations, and models used for our evaluation.
@@ 3.1 Data
Resumes are sourced through social media platforms (LinkedIn, Slack, X). Given the authors’ professional circles, the sample skews heavily toward tech and academic professionals. For privacy reasons, we anonymize resumes using Presidio to mask PII entities (Microsoft, ). To further mitigate privacy concerns and enable dataset release, we use the collected resumes as examples to generate synthetic resumes. We generate 525 resumes across 22 professions using Cohere’s Command-R model (Cohere, 2024). All synthetic resumes are free of explicit demographic information, until added during experimentation.
In addition, we use a publicly available resume dataset from Kaggle (Bhawal, 2021) to increase coverage and generalization. These resumes differ in two notable ways: (1) they are less structured and formatted than generated ones, and (2) they include a more diverse set of fields (e.g., construction, fitness, etc.). We sample 1175 Kaggle resumes across 24 fields. More details about dataset curation and statistics are provided in Appendix A.4 - A.6.
@@
Our resume dataset consists of two types: synthetic resumes generated for specific roles (e.g., data analyst) and actual resumes labeled with broader field categories (e.g., construction). For each profession/field (we choose 11 each from generated and Kaggle resumes), we carefully select 7 detailed LinkedIn job postings, resulting in 154 job postings.
@@ 3.2 Demographic Perturbations
We use names as a proxy for gender and racial information. All resumes are initially free of names; we add them using the curated set from Yin et al. (2024). [^5] Uses voter registration data from North Carolina to identify demographically-distinct names. We consider four demographic groups, each with 100 unique names: Black female (FB), White female (FW), Black male (MB), and White male (MW). Following Wilson and Caliskan (2024), we only vary the first name and fix “Williams” as the last name for all groups.
In actual resumes, demographic information can be encoded in more than just names. Therefore we perform an additional augmentation step that adds extracurricular information using Command-R [^6] Awards, clubs and leadership, and mentorship and volunteering experiences that are reflective of the individual’s background and identity (see Appendix A.6 and A.7). to the resumes (similar to Glazko et al. (2024)) Adding this information can reinforce demographic signal by providing both explicit and implicit cues.